OpenStudy (anonymous):

If a man earns $420 interest in one year at 7%, what was the amount of money invested?

OpenStudy (anonymous):

If a man earns $420 interest in one year at 7%, what was the amount of money invested? $29.40 $60.00 $600.00 $6,000.00

OpenStudy (anonymous):

420 = [principle](.07)^n , where n is the years, this case 1

OpenStudy (anonymous):

420/.07 = principle investment

OpenStudy (anonymous):

@Fazeelayaz is incorrect

OpenStudy (anonymous):

what would it be then???

OpenStudy (anonymous):

yes sory the formula is Sn=ar^n

OpenStudy (anonymous):

Sn is sum a is in this case principle investment and r is ur iinterst

OpenStudy (anonymous):

?

OpenStudy (anonymous):

if u have studied geometric serries then there is a formula if serries increases by addition of certain amount S=a*r^n

OpenStudy (anonymous):

oh ok soo what would the answer be to the question???

OpenStudy (raden):

let "x" is the amount of money invested, so it can be (7/100)*x = $420 x = $6,000.00

OpenStudy (anonymous):

How much pure water must be mixed with 10 liters of a 25% acid solution to reduce it to a 10% acid solution? 11 L 15 L 25 L

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

dilution problem: (10 L)(.025 M)=(X L)(.10 M) solve for X